OpenStudy (stormswan):

Can someone please help me? I really need help with this. Iwill give a medal. Find the area of the following shape. You must show all work to receive credit.

OpenStudy (hitaro9):

Think you can take a picture or draw the shape?

OpenStudy (hitaro9):

Okay. If you're ever given something that seems unnecesary complicated, always try to take it down to something very simple. If you were given a similar problem, but it was just a rectangle or a triangle on a similar grid, could you do it?

OpenStudy (stormswan):

uhh maybe. would it just be multiplying the sides?

OpenStudy (stormswan):

if that is the case then can't i just make that one shape in 2 shapes.. and find out the different areas?

OpenStudy (hitaro9):

Right. For a rectangle to get the area you multiple the side lengths. For a right triangle you multiple the side lengths and divide by two Does this picture help at all? http://puu.sh/idZon/229348e264.png

OpenStudy (hitaro9):

If you can find the area of each individual shape and add them up, you'll get the whole area

OpenStudy (stormswan):

So for the rectangle it would be 6x6=36 right?

OpenStudy (hitaro9):

Yep! 36 units squared

OpenStudy (stormswan):

and for the triangle i would do... 3x3 divided by 1/2? which would be 4.5?

OpenStudy (hitaro9):

Which triangle are you looking at? I think you might've miscounted the side length

OpenStudy (stormswan):

the one you drew.... of wait. it's 4 x 6. my bad. and then the answer would be 12?

OpenStudy (hitaro9):

Okay hold on

OpenStudy (hitaro9):

http://puu.sh/idZW3/9e91af3666.png

OpenStudy (hitaro9):

That triangle is 3 by 4. Does that make sense?

OpenStudy (stormswan):

ohhhhhhhh okay, nvm omg. I was doing it the wrong way, yeah. i get it. Then it's 6.

OpenStudy (stormswan):

so then the final answer is 42. okay

OpenStudy (stormswan):

Thank you!!!!!

OpenStudy (hitaro9):

Wait. There are two other triangles

OpenStudy (stormswan):

wait wut .-.

OpenStudy (hitaro9):

It's easy from here, but don't forget them haha

OpenStudy (hitaro9):

http://puu.sh/idZon/229348e264.png

OpenStudy (hitaro9):

Sometime's it's important to look back at the bigger picture. There are 3 triangles and one rectangle

OpenStudy (stormswan):

so... 3x4 div by 1/2= 6 3x6 div by 1/2= 9 6x6 div by 1/2= 18

OpenStudy (stormswan):

and for the rectangle it would be 6x6=36

OpenStudy (stormswan):

and the final answer is..........69 :) am I right?

OpenStudy (hitaro9):

Oh sorry. Umm, where are you seeinga 6x6 triangle?

OpenStudy (hitaro9):

Oh. Keep in mind that you only look at the bases of a right triangle for its area. So a triangle that looks like this http://puu.sh/ie16I/1549b72568.png Has bases 1x6, so it's 6x1 div by 2= 3

OpenStudy (hitaro9):

Other than that I think it looks good

OpenStudy (hitaro9):

Should be 54 rather than 69

OpenStudy (stormswan):

i thought it was 56? not 54...

OpenStudy (hitaro9):

Oh maybe I added wrong 1 sec

OpenStudy (hitaro9):

36+6+9+3 36+6 = 42 42+9 = 51 51+3=54

OpenStudy (hitaro9):

If I'm making some dumb arithmetic mistake feel free to point it out

OpenStudy (stormswan):

ohhh okay, sorry nvm

OpenStudy (stormswan):

Thank you for your help!!

OpenStudy (hitaro9):

Okay. Any questions about what we did? Does each step make sense?

OpenStudy (anonymous):

Sum the area of triangle ABF and trapezoid EBCD. Refer to the attachment.